Homework: Chapter 4 Homework Save < 13 of 15 (11 complete) HW Score: 73.33%, 11 of...


Café Michigan's manager, Gary Stark, suspects that demand for mocha latte coffees depends on the price being charged. Based on historical observations, Gary has gathered the following data, which show the numbers of these coffees sold over six different price values: 

PriceNumber Sold
$2.70770
$3.40510
$2.10980
$4.20240
$3.10315
$4.00490

Using simple linear regression and given that the price per cup is $1.85, the forecasted demand for mocha latte coffees will be _______  cups (enter your response rounded to one decimal place).
Answer #1

the trendline equation is y = -295.3x + 1510.5

Explanation:

  • linear trend forecast- select the period and demand and insert an X-Y chart. After making X-Y chart in Excel. there is an option to show the trendline. After that, there are>more options>show the trend equation on the chart. To find the forecast, replace the value of x in the equation with a period no. and find the value of Y.

Answer: For forecast of price $1.85, keep x= 1.85

F-1.85 = y = -295.3*1.85 + 1510.5= 964.2